# Herding Unmasked: Insights into Cryptocurrencies, Stocks and US ETFs

### v1

@@ 3.1 Data Sources
Each time series covers the period from 01/04/2019 until 03/05/2023 and has a resolution of 30 minutes. Since the official trading time of traditional markets is between 09:30 a.m. and 4:00 p.m., we select this timeframe as the boundary for our time series, meaning that the cryptocurrency data is only available within this trading time, even though these coins are traded nonstop. Other price values outside of this timeframe are not considered in our experiments. This practice facilitates the comparison between assets from different types of investment vehicle and also preserves their equality. A list of assets in each type of investment vehicle is given in Appendix B.
@@ 3.2 Financial Sector Classification
We note that the number of assets in each stock sector is different and depends on its asset’s market capitalization. For instance, Technology, Finance and Healthcare stocks tend to have very high market capitalization while Communication Services and Real Estate tend to have lower market capitalization [97]. Consequently, among the 11 stock sectors, Utilities, Materials, Real Estate and Communication Services exhibited a limited number of entities. Therefore, this study concentrates only on the remaining 7 sectors. Further details for this classification are provided in Appendix C.
@@ 3.4 Louvain Community Detection
Community detection aims to partition nodes of a graph into separate groups (called communities) by maximizing the similarity between nodes within a community while maximizing the distinction between different communities. In other words, this detection results in communities where nodes belonging to a community are most similar to each other while nodes from different communities exhibit different characteristics. For this technique, Louvain community detection algorithm is well-known for its applications in large networks [30]. Additionally, this approach is rather straightforward and is also adopted widely in various research topics [103, 104, 105]. Since our graph has more than 200 nodes, this method is unquestionably well-suited for this case.
@@
From this idea, the process of Louvain community detection comprises two phases, as shown in Pseudocode 1. Moreover, Fig 1 illustrates the process of constructing communities from a set of time series.
@@
Figure 1: The process of implementing community detection from a set of time series. For more information, see Clarification 1..
@@ 3.5 Herding Detection Strategies
Regarding Eq 3, a herding exists when the parameter $\beta_{2}$ is significantly negative (a negative coefficient with a confidence level of at least 90%). Intuitively, we expect this coefficient to be negative because it follows the idea of the herding behavior, i.e. a dramatic increase or decrease in the market return ($r_{m,t}^{2}$) coincides with a decline (negative $\beta_{2}$) in the dispersion between corresponding assets’ returns ($CSAD_{t}$). Eq 4 refers to a more detailed version of Eq 3 in which we separate the market condition into up (i.e. positive market return) and down (i.e. non-positive market return) so that the herding can be detected during the up and down market condition separately, instead of considering all market conditions (i.e. both up and down equally) as declared in the first form. The use of both forms for herding detection comes from an expectation that there might be periods when herding does not exist when considering both up and down market conditions at the same time but it might be found in the up or down market condition only. To this end, the same logic of Eq 3 is applied to Eq 4 wherein a significantly negative $\gamma_{2}$ refers to a herding signal during the up market condition while a significantly negative $\gamma_{3}$ refers to a herding signal during the down market condition.
@@ 3.6 An Early Experiment: Herding Behavior throughout the Entire Period
Table 1 provides the estimation of $\beta_{2}$ from Eq 3 and $\gamma_{2},\gamma_{3}$ from Eq 4 for each type of investment vehicle. All of them reveal signs of herding behavior over the considered timeline from April 2019 until May 2023 as the coefficient $\beta_{2}$ is statistically negative in all cases, even though their herding magnitudes are different. When distinguishing the market condition into up and down, herding is pronounced in both conditions for stocks and US ETFs, however, only the down condition in the cryptocurrency market exhibits this phenomenon.
@@
The parameters $\gamma_{2}$ and $\gamma_{3}$ are estimated from the equation:
@@
Fig 2 illustrates herding detection results and sector distribution of each community (as a result of applying the Louvain community detection algorithm) in the combination of all three types of assets over the entire period. Specifically, the number of communities found by the Louvain algorithm and the proportion of each sector in each community is illustrated in the lower sub-figure. Sectors are listed on the y-axis, each represented by a unique color. For each community, the longer the bar of a corresponding sector, the greater the sector’s percentage within that community. In the meantime, the upper sub-figure reveals the corresponding herding detection results for each community in the lower sub-figure. The grey bar means the herding based on CSAD model (Eq 3) is significant, the green bar means the herding based on CSAD for the up market condition (Eq 4) is significant, the red bar means the herding based on CSAD for down market condition (Eq 4) is significant. The higher the bar, the more pronounced the herding. The assets in each community are listed in Appendix D. We find that most communities reveal herding except for the two communities 7 and 8. Overall, it is shown that performing herding detection on the whole period tends to finish with a ’Yes’ result. However, it is almost impossible to conclude anything given this result since the examined period is long, leading to a lack of detail. Furthermore, all types of investment vehicle exhibit herding behavior so no unique patterns can be found. This ambiguity encourages us to split the original period into sub-periods so that the herding behavior can be examined on a finer time interval, which enhances the detail level of the experiment. The information for this timeline division is proposed in the next section.
(a) Herding Detection Results
(b) Sector Distribution
Figure 2: Herding detection on each community (a) and its sector distribution (b) over the entire period from 01/04/2019 until 03/05/2023. More information on each sub-figure can be seen at Clarification 2..
@@ 3.7 Timeline Division
Notably, after Covid-19’s first wave, the financial markets began to enter the recovery time and most assets in both traditional and cryptocurrency markets have witnessed a surge in prices since the beginning of 2021. This was a period when most assets reached their all-time highest values. Since 2022, the markets have calmed down gradually, marking the end of the bullish time in April 2022 for most assets. Thus, we also include this Bull market time in our time division. Table 2 displays an explicit division for the original period.
Table 2: Five sub-periods declaration.
Name Pre-Covid-19 Covid-19 pandemic Post-Covid-19 Bull market
Ukraine-Russia
conflict
Time interval
01/04/2019
to
31/12/2019
01/01/2020
to
30/06/2020
01/07/2020
to
31/12/2020
01/01/2021
to
30/04/2022
01/05/2022
to
03/05/2023
@@ 4.1 Herding Detection on Each Type of Investment Vehicle
In this part, we tackle the first research question by uncovering herding signals in each type of investment vehicle across the five aforementioned sub-periods. Subsequently, answering various sub-questions such as: Which events cause the phenomenon? How strong is it? How are investors’ reactions linked to the herding? How resilient is the herding? Then, a comparison between the three types of investment vehicle is conducted.
Table 3 provides the estimation of $\beta_{2}$ from the CSAD model (without distinguishing up and down market conditions) while Table 4 provides the estimation of $\gamma_{2}$, and $\gamma_{3}$ from CSAD for the up market condition and CSAD for the down market condition, respectively, for each type of investment vehicle. Only values with the star symbol are statistically significant. Herd behavior is revealed at a period only when its corresponding coefficient (either $\beta_{2}$, $\gamma_{2}$ or $\gamma_{3}$) is statistically significantly negative. The more negative the value, the more pronounced the herding.
@@
Pre-Covid-19 0.0801* -4.550* -0.0894*
Covid-19 pandemic -0.0416* -0.995* -0.0281*
Post-Covid-19 -0.0524* 8.684* 0.1032*
Bull market -0.0271* 31.398* 0.0265
U-R conflict 0.0561* 0.155 0.0144
@@
Up Market Condition ($\gamma_{2}$) Down Market Condition ($\gamma_{3}$)
@@
Pre-Covid-19 0.1133* -2.064 0.0001 0.0585* -5.085* -0.0951*
Covid-19 pandemic -0.0379* -1.297* -0.0305* -$0$998* -0.0270*
Post-Covid-19 -0.0471* 9.528* 0.0774* -0.0474* -0.330 0.0988*
Bull market 0.0229* 6.725* 0.0603** -0.0276* 33.658* 0.0262
U-R conflict 0.0511* 0.464 0.0198 0.0604* -0.417 0.0049
The parameters $\gamma_{2}$ and $\gamma_{3}$ are estimated from the equation
@@
Starting with the cryptocurrency market, the herding coefficient $\beta_{2}$ reveals herding signals in three sub-periods, including during the Covid-19 pandemic, the Post-Covid-19 and the Bull market time. Surprisingly, the strongest herding is found in the Post-Covid-19 while the peak of the Covid-19 pandemic, which we expected to show the strongest herding behavior, only comes in second place. On the other hand, the Pre-Covid-19 and Ukraine-Russia conflict sub-periods disclose an absence of herding behavior. The results remain mostly unchanged after separating the market into up and down conditions. Specifically, the coefficients $\gamma_{2}$ and $\gamma_{3}$ are only statistically significantly negative during the sub-periods when $\beta_{2}$ is negative, indicating that either the up or down market condition or both conditions reveal the evidence of herding in these stages. However, while herding is pronounced in both up and down market conditions during the Covid-19 pandemic and Post-Covid-19 sub-periods, only the downturn times in the Bull market time have the appearance of herding. These results sufficiently capture the prevailing sentiment among cryptocurrency investors. Specifically, they seem to be prone to be influenced by market fluctuations and breaking news on the internet [107], leading to similar investment decisions, i.e. selling off their investments when the market experiences a downturn or bad news circulates on the internet to avoid a loss while purchasing new investments during bullish phases or when positive news about cryptocurrencies spreads among the general public due to the Fear of Missing Out (FOMO) on an increase in the price of the coins[90]. The high degree of sensitivity of investors in this digital market may be attributed to the dominance of irrational (naïve) participants, as confirmed in different studies [8], these people participate in the cryptocurrency market primarily because of its great potential for huge profit despite their lack of investment experience. Therefore, they tend to ignore their own analysis and information by following the decisions made by the public due to their lack of financial knowledge [27]. A greater herding magnitude during the Post-Covid-19 compared to the peak of the pandemic can be attributed to the cryptocurrency market’s tendency to be more influenced by positive circumstances than negative ones [108, 19]. During this time, the market started to recover from the decline caused by the pandemic, experiencing a price surge since the fourth quarter of 2020. An upward trend in prices along with positive word of mouth, as well as financial news, triggered the positive sentiment of investors together with the FOMO phenomenon that eventually caused a purchase herding in the market [27, 96, 21]. When it comes to the Bull market time, although the herding is evident, it appears to be weaker than other sub-periods. This might be because many investors had already purchased the coins at the earlier stage and were more likely to hold onto those coins during this time because of the positive expectations of an upward trend in the future. Thus, they barely purchased at this time, instead, the herding might come from them executing a sell action. Indeed, we notice that there was a downturn phase in this sub-period when the market significantly shrank. Hence, it is possible that a portion of investors began selling off their investments as they either felt panicked by the price drop or were motivated to take profits, resulting in a period of herding [90, 106, 109, 90]. Our result when examining the cryptocurrency market in up and down conditions also shows that only $\gamma_{3}$ is negative during this phase, meaning that only the downturn time in the Bull market time reveals the herding behavior, which reinforces our hypothesis. Overall, our results of herding detection are in line with existing literature which also demonstrates evidence of herding during the first outbreak of the pandemic/ 2020 global economic crisis and the periods of bullish markets, while there was no herding signal found before the pandemic and after the bullish time. [7, 21, 96, 110].
@@
The herding behavior observed in the stock and US ETF markets displays similar patterns, with only the first two sub-periods, namely Pre-Covid-19 and Covid-19 pandemic, having signs of herding behavior. These time intervals cover critical events that significantly impacted traditional markets, i.e. the US-China trade war during the year 2019 and the severe hit of the contagious health disease Covid-19 on the global economy and financial markets in the first half of 2020. Notably, the US-China trade war took place throughout 2019 and continually affected the economies of not only the US and China but also other countries and continents [45, 39]. Moreover, investors have also been bombarded by both negative and positive news about this event, with the largest contribution coming from President Trump who often used social networks to post updates on the trade war. Consequently, such news dragged the markets up and down frequently [42]. Indeed, it has been shown that there was a selling wave of assets during this year [43]. This explains the more pronounced herding within the negative market return compared to the positive market return during the Pre-Covid-19, in both stock and US ETF markets.
@@
The herding in the stock and US ETF markets delivers two interesting results. Firstly, herding is more pronounced during the Pre-Covid-19 compared to the Covid-19 pandemic sub-period. Therefore, although the pandemic led to greater losses and fluctuations in traditional financial markets, it is noteworthy that investors exhibited more intense herding behavior before the pandemic, particularly during the turbulent time caused by the US-China trade war. One possible reason for this is the constant fluctuations of markets throughout 2019 when the trade war became more and more intense. In particular, different tariffs and quotas were imposed between these two countries throughout this year, directly impacting financial markets worldwide. As a result, the constant release of negative and positive news gave rise to frequent market movements. Thus, investors actively reacted to market situations, causing herding behavior on a regular basis. The impact of the Covid-19 pandemic was truly severe and damaging within the second quarter of 2020 but this was followed by a gradual recovery. Hence, investors panicked during the peak of the pandemic but things quickly got back to normal, and as they started to make objective investment decisions again, herding then lessened. Another possible reason is that investors in traditional markets are more influenced by macroeconomic factors as well as policy uncertainty. Specifically, US interest rates increased because the government expected the economy to grow in 2019. Furthermore, the unemployment rate increased due to the trade war and GDP in 2019 also fell to its lowest since the 2008 crisis [111, 112, 4]. These macroeconomic factors had the potential to scare traditional market traders.
Secondly, the magnitude of herding behavior in the market composed of US ETF appears to dramatically surpass that observed in the stock market (and also the cryptocurrency market). We explain this distinction by examining the CAPM beta score [113] of each asset. An asset’s beta score measures its returns’ tendency to follow the overall market. In other words, it measures how likely an asset’s returns increase (or decrease) when the overall market’s returns increase (or decrease) where the overall market tends to be the S&\&&P 500 index. A beta of unity means that the asset’s movements are exactly the same as the overall market while less than unity indicates weaker movements, resulting in a lower return compared to the overall market and greater than unity indicates stronger movements, resulting in a higher return compared to the overall market. From Fig 3, we find that beta scores of US ETFs tend to be close to each other and show a tight range over those of the overall market. On the other hand, the beta scores of stocks show a wider dispersion. However, all of these traditional assets exhibit a correlation with the overall market to some extent as their beta scores are close to unity. By contrast, cryptocurrencies exhibit a different behavior since not all of them follow the movements of the S&\&&P 500 index, indicated by the beta scores of almost zero. Instead, cryptocurrencies tend to follow the movements of Bitcoin. Additionally, to compare the beta scores between stocks, US ETFs and cryptocurrencies numerically, we also calculate the average distance of assets’ beta scores in each type of investment vehicle compared to the overall market as measured by the S&\&&P500. We find that the beta scores of US ETFs are closer to unity, as compared to stocks and cryptocurrencies which reveal a higher distance value compared to US ETFs. This result is shown in Table 5. Therefore, US ETFs have a higher tendency to follow the movements of the overall market (S&\&&P500) when compared to underlying stocks and cryptocurrencies, which explains the high magnitude of the herding behavior of US ETFs. Another possible reason might be the high liquidity of these stock ETFs. Specifically, because the liquidity of ETFs is higher than stocks [114, 115], their prices mirror investors’ reactions (e.g. buy or sell) quicker [116]. Consequently, the chance for their price movements to become alike is higher than that of underlying stocks, resulting in stronger herding. On the other hand, stocks tend to have lower liquidity [114, 115] and the degree of liquidity across different stocks might be very different. Thus, even if investors react similarly to different stocks at the same time, the difference in different stocks’ liquidity causes a lag between their price movements and lessens the herding.
Figure 3: Box Plots of Beta Scores in each Market. The overall market is the S&\&&P 500 index for traditional assets (i.e. stocks and ETFs) while both S&\&&P 500 index and Bitcoin are used as the overall market for cryptocurrencies.
Table 5: Average distance between assets’ beta scores and the market index (S&\&&P500 for traditional assets and Bitcoin for cryptocurrencies) in each market. The smaller the distance is, the more similar the assets are to the market index.
Metrics
Stocks
(S&\&&P500 base)
US ETFs
(S&\&&P500 base)
Cryptocurrencies
(Bitcoin base)
MAE 0.272 0.156 0.178
RMSE 0.219 0.124 0.158

### v2

@@ 3.1 Data Sources
Each time series covers the period from 01/04/2019 until 03/05/2023 and has a resolution of 30 minutes. Since the official trading time of traditional markets is between 09:30 a.m. and 4:00 p.m., we select this timeframe as the boundary for our time series, meaning that the cryptocurrency data is only available within this trading time, even though these coins are traded nonstop. Other price values outside of this timeframe are not considered in our experiments. This practice facilitates the comparison between assets from different classes and also preserves their equality. A list of assets in each class is given in S1 Table..
@@ 3.2 Financial Sector Classification
We note that the number of assets in each stock sector is different and depends on its asset’s market capitalization. For instance, Technology, Financials and Healthcare stocks tend to have very high market capitalization while Communication Services and Real Estate tend to have lower market capitalization [110]. Consequently, among the 11 stock sectors, Utilities, Materials, Real Estate and Communication Services exhibited a limited number of entities. Therefore, this study concentrates only on the remaining 7 sectors. Further details for this classification are provided in S2 Table..
@@ 3.4 Louvain Community Detection
Community detection aims to partition nodes of a graph into separate groups (called communities) by maximizing the similarity between nodes within a community while maximizing the distinction between different communities. In other words, this detection results in communities where nodes belonging to a community are most similar to each other while nodes from different communities exhibit different characteristics. In the context of financial markets, where nodes represent financial assets, a community in this case comprises assets that have most similar price movements over a certain period of time. On the other hand, assets from different communities exhibit less similar price movement patterns. For this technique, Louvain community detection algorithm is well-known for its applications in large networks [31]. Additionally, this approach is rather straightforward and is also adopted widely in various research topics [116, 117, 118, 119]. Since our graph has more than 200 nodes, this method is unquestionably well-suited for this case.
@@
From this idea, Louvain community detection algorithm comprises two phases, as shown in Pseudocode 1. Furthermore, the whole process of constructing communities from a set of financial time series is illustrated and interpreted in S1 Appendix..
@@ 3.5 Herding Detection Strategies
Regarding Eq 3, a herding exists when the parameter $\beta_{2}$ is significantly negative (a negative coefficient with a confidence level of at least 90%). Intuitively, we expect this coefficient to be negative because it follows the idea of the herding behavior, i.e. a dramatic increase or decrease in the market return ($r_{m,t}^{2}$) coincides with a decline (negative $\beta_{2}$) in the dispersion between corresponding assets’ returns ($CSAD_{t}$). Eq 4 refers to a more detailed version of Eq 3 in which we separate the market condition into up (i.e. positive market return) and down (i.e. non-positive market return) so that the herding can be detected during the up and down market condition separately, instead of considering all market conditions (i.e. both up and down equally) as declared in the first form. The use of both forms for herding detection comes from an expectation that there might be periods when herding does not exist when considering both up and down market conditions at the same time but it might be found in the up or down market condition only. To this end, the same logic of Eq 3 is applied to Eq 4 wherein a significantly negative $\gamma_{2}$ refers to a herding signal during the up market condition while a significantly negative $\gamma_{4}$ refers to a herding signal during the down market condition.
3.6 Contagion Detection Strategy
To capture the contagion effects between multiple time series, the Vector Autoregression (VAR) model has emerged as a foundational econometric tool. The VAR model, first introduced by Sims [121], is a statistical model that captures the linear interdependencies among multiple time series by allowing each variable to be regressed on lagged values of itself and other variables in the system [122]. Therefore, this model can identify bidirectional shock transmission between any two time series, capturing shocks moving from one series to the other and vice versa. Specifically, given a set of $k$ time series $\{y_{1},y_{2},...,y_{k}\}$, the VAR model of order $p$ to assess the contagion effects between these $k$ time series can be expressed as:
$\begin{bmatrix}y_{1,t}\\ y_{2,t}\\ \vdots\\ y_{k,t}\end{bmatrix}=\begin{bmatrix}c_{1}\\ c_{2}\\ \vdots\\ c_{k}\end{bmatrix}+\begin{bmatrix}a_{11}^{t-1}&\cdots&a_{1k}^{t-1}\\ a_{21}^{t-1}&\cdots&a_{2k}^{t-1}\\ &\vdots&\\ a_{k1}^{t-1}&\cdots&a_{kk}^{t-1}\end{bmatrix}\begin{bmatrix}y_{1,t-1}\\ y_{2,t-1}\\ \vdots\\ y_{k,t-1}\end{bmatrix}+\dots+\begin{bmatrix}a_{11}^{t-p}&\cdots&a_{1k}^{t-p}\\ a_{21}^{t-p}&\cdots&a_{2k}^{t-p}\\ &\vdots&\\ a_{k1}^{t-p}&\cdots&a_{kk}^{t-p}\end{bmatrix}\begin{bmatrix}y_{1,t-p}\\ y_{2,t-p}\\ \vdots\\ y_{k,t-p}\end{bmatrix}+\begin{bmatrix}\varepsilon_{1}\\ \varepsilon_{2}\\ \vdots\\ \varepsilon_{k}\end{bmatrix}$ (5)
[…]
@@ 3.7 An Early Experiment: Herding Behavior throughout the Entire Period
Table 1 provides the estimation of $\beta_{2}$ from Eq 3 and $\gamma_{2},\gamma_{4}$ from Eq 4 for each type of investment vehicle. All of them reveal signs of herding behavior over the considered timeline from April 2019 until May 2023 as the coefficient $\beta_{2}$ is statistically negative in all cases, even though their herding magnitudes are different. When distinguishing the market condition into up and down, herding is pronounced in both conditions for stocks and US ETFs, however, only the down condition in the cryptocurrency market exhibits this phenomenon.
@@
The parameters $\gamma_{2}$ and $\gamma_{4}$ are estimated from the equation:
@@
Fig 1 illustrates herding detection results and sector distribution of each community (as a result of applying the Louvain community detection algorithm) in the combination of all three types of assets over the entire period. Specifically, the number of communities found by the Louvain algorithm and the proportion of each sector in each community is illustrated in the lower sub-figure (Fig 1b). For each community, the dominance of each financial sector within this community is indicated by the length of its corresponding horizontal bar. Sectors are listed on the y-axis, each represented by a unique color, the longer the bar of a corresponding sector, the greater the sector’s percentage within that community. In the meantime, the upper sub-figure (Fig 1a) reveals the corresponding herding detection results for each community in the lower sub-figure (Fig 1b). To visualize these herding coefficients, we use their absolute values and refer to them as herding magnitudes, if there is no evidence of herding, the herding magnitude is equal to 0. In this regard, the grey bar means the herding based on CSAD model (Eq 3) is significant, the green bar means the herding based on CSAD for the up market condition (Eq 4) is significant, the red bar means the herding based on CSAD for down market condition (Eq 4) is significant. The higher the bar, the more pronounced the herding. The assets in each community are listed in S3 Table.. Overall, it is shown that performing herding detection on the whole period tends to finish with a ’Yes’ result. However, it is almost impossible to conclude anything given this result since the examined period is long, leading to a lack of detail. Furthermore, all types of investment vehicle exhibit herding behavior so no unique patterns can be found. This ambiguity encourages us to split the original period into sub-periods so that the herding behavior can be examined on a finer time interval, which enhances the detail level of the experiment. The information for this timeline division is proposed in the next section.
Figure 1: Herding detection on each community (a) and its sector distribution (b) over the entire period from 01/04/2019 until 03/05/2023. More information on each sub-figure can be seen at S1 Text..
@@ 3.8 Timeline Division
Notably, after Covid-19’s first wave, the financial markets began to enter the recovery time and most assets in both traditional and cryptocurrency markets have witnessed a surge in prices since then. This was a period when most assets reached their all-time highest values. Since 2022, the markets have calmed down gradually, marking the end of the bullish time for most assets [123]. Thus, we also consider this bull market time in our timeline division.
To this end, we partition the original timeline into 7 sub-periods as shown in Table 2. We provide some explanations for this division of sub-periods in this table. Firstly, we choose the date 30/06/2020 as a cutoff for the Covid-19 pandemic sub-period because the financial markets have recovered to their Pre-Covid-19 performance and the general public attention towards the pandemic has been lessened significantly in July 2020 [59]. Since then, both cryptocurrency and traditional markets continued surging in prices until the end of 2021. In other words, the financial markets entered bull market conditions after overcoming the severe hit by the global economic crisis in March 2020. Ideally, the period from 01/07/2020 to 23/02/2023 (before the Ukraine-Russia conflict onset) can be called a bull market time, but this period is too long. Therefore, we divide it into 3 sub-periods. Lastly, the remaining time coincides with the Ukraine-Russia conflict event. However, this period is also too long compared to the previous ones. Therefore, we divide it into 2 sub-periods.
Table 2: Timeline division in this study. The original period (01/04/2019 - 03/05/2023) is segmented into 7 sub-periods based on key historical events and market conditions
Sub-period Name Time Interval
1 Pre-Covid-19 01/04/2019 to 31/12/2019
2 Covid-19 Pandemic 01/01/2020 to 30/06/2020
3 Bull Time 1 01/07/2020 to 31/01/2021
4 Bull Time 2 01/02/2021 to 31/08/2021
5 Bull Time 3 01/09/2021 to 23/02/2022
6 Ukraine-Russia Conflict 1 24/02/2022 to 30/09/2022
7 Ukraine-Russia Conflict 2 01/10/2022 to 03/05/2023
@@ 4.1 Herding Detection on Each Type of Investment Vehicle
In this part, we tackle the first research question by uncovering herding signals in each type of investment vehicle across the seven aforementioned sub-periods. Subsequently, answering various sub-questions such as: Which events cause the phenomenon? How strong is it? How are investors’ reactions linked to the herding? How resilient is the herding? Then, a comparison between the three types of investment vehicle is conducted.
Table 3 provides the estimation of $\beta_{2}$ from the CSAD model (without distinguishing up and down market conditions) while Table 4 provides the estimation of $\gamma_{2}$, and $\gamma_{4}$ from CSAD for the up market condition and CSAD for the down market condition, respectively, for each type of investment vehicle. Only values with the star symbol are statistically significant. Herd behavior is revealed at a period only when its corresponding coefficient (either $\beta_{2}$, $\gamma_{2}$ or $\gamma_{4}$) is statistically significantly negative. The more negative the value, the more pronounced the herding.
@@
Pre-Covid-19 6.3223* -4.5498* -9.1788*
Covid-19 Pandemic -4.7204* -0.9952* -2.7261*
Bull Time 1 -6.7795* 8.1034* 9.2020*
Bull Time 2 -3.7911* 38.3803* -10.1002**
Bull Time 3 1.0469* 1.5452 3.8338
U-R Conflict 1 0.4652** 1.6437** 5.9794*
U-R Conflict 2 -9.1943* -0.4417 -0.5448
@@
Up Market Condition ($\gamma_{2}$) Down Market Condition ($\gamma_{4}$)
@@
Pre-Covid-19 9.1014* -2.0638 -1.6660 4.5191* -5.0853* -10.0953*
Covid-19 Pandemic -4.3471* -1.2972* -3.2433* -4.6894* -0.9982* -2.5910*
Bull Time 1 -6.4200* 8.9108* 7.0201* -6.1360* -0.5774 6.5425**
Bull Market 2 0.4969 10.9087* -3.3375 -3.9754* 38.9335* -14.5454*
Bull Time 3 2.2923* -5.5150** -0.5228 1.0751* 3.5424* 6.8038*
U-R Conflict 1 0.0358 7.3115* 15.5605* 0.6319* 0.6063 3.7276***
U-R Conflict 2 -11.9781* -0.2959 -0.7646 -7.8995* -0.5792 -0.3428
The parameters $\gamma_{2}$ and $\gamma_{4}$ are estimated from the equation
@@
Starting with the cryptocurrency market, the herding coefficient $\beta_{2}$ reveals herding signals in four sub-periods, including during the Covid-19 Pandemic, the first and second Bull Time, and the second Ukraine-Russia Conflict. Surprisingly, the strongest herding signals are found in the first Bull Time and the second Ukraine-Russia Conflict sub-periods while the peak of the Covid-19 pandemic, which we expected to show the strongest herding behavior, only comes in third place. On the other hand, the Pre-Covid-19, the third Bull Time and the first Ukraine-Russia Conflict sub-periods disclose an absence of herding behavior. The results remain mostly unchanged after separating the market into up and down conditions. Specifically, the coefficients $\gamma_{2}$ and $\gamma_{4}$ are only statistically significantly negative during the sub-periods when $\beta_{2}$ is negative, indicating that either the up or down market condition or both conditions reveal the evidence of herding in these stages. However, while herding is pronounced in both up and down market conditions in most cases, only the downturn times in the second sub-period where Bull Market conditions prevail have the appearance of herding. These results sufficiently capture the prevailing sentiment among cryptocurrency investors. Specifically, they seem to be influenced by market fluctuations and breaking news[124], leading to similar investment decisions, i.e. selling off their investments when the market experiences a downturn or bad news circulates on the internet to avoid a loss while purchasing new investments during bullish phases or when positive news about cryptocurrencies spreads among the general public due to the Fear of Missing Out (FOMO) on an increase in the price of the coins[92]. The high degree of sensitivity of investors in this digital market may be attributed to the dominance of irrational (naïve) participants, as confirmed in different studies [8], these people participate in the cryptocurrency market primarily because of its great potential for huge profit despite their lack of investment experience. Therefore, they tend to ignore their own analysis and information by following the decisions made by the public due to their lack of financial knowledge [27]. A greater herding magnitude during the first Bull Time and the second Ukraine-Russia Conflict sub-periods compared to the peak of the pandemic (Covid-19 Pandemic sub-period) can be attributed to the cryptocurrency market’s tendency to be more influenced by positive circumstances than negative ones [125, 19]. Specifically, during the first Bull Time, the market started to recover from the decline caused by the pandemic, experiencing a price surge since then. An upward trend in prices along with positive word of mouth, as well as financial news, triggered the positive sentiment of investors together with the FOMO phenomenon that eventually caused a purchase herding in the market [27, 97, 21]. When it comes to the second Bull Time, although the herding is evident, it appears to be weaker than other sub-periods. This might be because many investors had already purchased the coins at the earlier stage and were more likely to hold onto those coins during this time because of the positive expectations of an upward trend in the future. Thus, they barely purchased at this time, instead, the herding might come from them executing a sell action. Indeed, we notice that there was a downturn phase in this sub-period when the market significantly shrank. Hence, it is possible that a portion of investors began selling off their investments as they either felt panicked by the price drop or were motivated to take profits, resulting in a period of herding [92, 120, 126, 92]. Our result when examining the cryptocurrency market in up and down conditions also shows that only $\gamma_{4}$ is negative during the second Bull Time, meaning that herding behavior is much more intense during the downturn time compared to the upturn one, which reinforces our hypothesis. Moreover, the same circumstance is also seen in the second Ukraine-Russia Conflict sub-period. During this time, financial markets experienced an upward trend and cryptocurrencies are no exception to this. Specifically, herding behavior in the cryptocurrency market is found in both up and down market conditions with the herding being more intense during times of positive market returns. Overall, results of herding detection among cryptocurrencies market are in line with existing literature which also demonstrates evidence of herding during the first outbreak of the pandemic/ 2020 global economic crisis and the periods of bullish markets [7, 21, 97, 127].
@@
The herding behavior observed in the stock and US ETF markets displays similar patterns most of the time, with the first two sub-periods, namely Pre-Covid-19 and Covid-19 Pandemic, having signs of herding behavior while most remaining sub-periods show no evidence of herding.These first two sub-periods cover critical events that significantly impacted traditional markets, i.e. the US-China trade war during the year 2019 and the severe hit of the contagious health disease Covid-19 on the global economy and financial markets in the first half of 2020. Notably, in the Pre-Covid-19 period (Tables 3 and 4), herding behavior in both stocks and US ETFs was significantly stronger during times of negative market returns than positive ones. This greater herding magnitude can be attributed to the tension of the US-China trade war throughout 2019, affecting not only the US and Chinese economies but also other countries and continents [46, 40]. Additionally, investors were continuously exposed to negative news about the trade war throughout the year [43]. Consequently, these circumstances led to a herding phenomenon among investors, especially during periods of declining asset prices. In fact, it has been shown that there was a selling wave of assets during this year [44]. The only exception was that stocks also exhibited a signal of herding behavior during the second Bull Time while US ETFs revealed evidence of this behavior during the third Bull Time. This herding behavior can be attributed to the continual surge in prices of the assets during these sub-periods [123], which caused similar purchasing and selling actions among investors with the aim of making the most profits.
@@
The herding in the stock and US ETF markets delivers two interesting results. Firstly, herding is more pronounced during the Pre-Covid-19 compared to the Covid-19 Pandemic sub-period. Therefore, although the pandemic led to greater losses and fluctuations in traditional financial markets, it is noteworthy that investors exhibited more intense herding behavior before the pandemic, particularly during the turbulent time caused by the US-China trade war. One possible reason for this is the constant fluctuations of markets throughout 2019 when the trade war became more and more intense. In particular, different tariffs and quotas were imposed between these two countries throughout this year, directly impacting financial markets worldwide. As a result, the constant release of negative and positive news gave rise to frequent market movements. Thus, investors actively reacted to market situations, causing herding behavior on a regular basis. The impact of the Covid-19 pandemic was truly severe and damaging within the second quarter of 2020 but this was followed by a gradual recovery. Hence, investors panicked during the peak of the pandemic but things quickly got back to normal, and as they started to make objective investment decisions again, herding then lessened. Another possible reason is that investors in traditional markets are more influenced by macroeconomic factors as well as policy uncertainty. Specifically, US interest rates increased in response to rising inflation. Furthermore, the unemployment rate increased due to the trade war and GDP in 2019 also fell to its lowest since the 2008 crisis [128, 129, 4]. These macroeconomic factors had the potential to scare traditional market traders.
Secondly, the magnitude of herding behavior in the market composed of US ETFs is generally weaker than in the stock and cryptocurrency markets, as indicated by the lower herding coefficients ($\beta_{2},\gamma_{2},\gamma_{4}$) for US ETFs compared to those for stocks and cryptocurrencies across most sub-periods. This distinction is possibly driven by the stylized facts of US ETFs: the US market itself has a well-established economic environment and a high proportion of rational investors [18, 22]. As a result, since investors are more inclined to rely on their own analysis and expectations to make investment decisions rather than mimicking other investors, the herding behavior is therefore reduced. This large number of rational investors may be a crucial factor in limiting herding behavior, since the herding frequency in stocks and US ETFs is lower than in cryptocurrencies, as evidenced by Tables 3 and 4. On the other hand, another stylized fact about ETFs is their low volatility (i.e. low risk) compared to other types of investment vehicle, which limits the potential for short-term profits. Therefore, investors tend to herd in other - more risky - types of investment vehicle to gain high profits while ETFs are preferable for long-term investments [130]. Remarkably, also in US ETFs, Tables 3 and 4 display an intense dispersion in price movements between these assets during the second Bull Time sub-period, with $\beta_{2}$ exceeding 38.0 and $\gamma_{4}$ nearing 39.0. This indicates a substantial divergence in investment decisions among investors in US ETFs during this time.